11.Evaluate the following Java expressions, assuming x and y are both declared as integer variables and x = 17and y = 6:


x/y+3     =                                         

x%y*4     =                                         

x-y/3     =                                       


1
Expert's answer
2021-07-13T06:38:19-0400
Answer:

a) x / y + 3 = 17 / 6 + 3 = 2 + 3 = 5 
b) x % y * 4 = 17 % 6 * 4 = 5 * 4 = 20 
c) x –y / 3=  17 - 6 / 3  = 15